integrate the following integral
∫ (2x – 3)^5 + sqrt(3x +2) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int [(2x-3)^5+\sqrt{3x+2}]dx
I = \int (2x-3)^5dx+\int \sqrt{3x+2}dx
I = \frac{(2x-3)^6}{12}+\int \sqrt{3x+2}dx
I = \frac{(2x-3)^6}{12}+\frac{2(3x+2)^{3/2}}{3.3}+c
I = \frac{(2x-3)^6}{12}+\frac{2(3x+2)^{3/2}}{9}+c